# A hidden Markov model for statistical arbitrage in international crude oil futures markets

### v2

@@ Abstract.
We study statistical arbitrage strategies in international crude oil futures markets. We analyse strategies that extend classical pairs trading strategies, considering two benchmark crude oil futures (Brent and WTI) together with the recently introduced Shanghai crude oil futures. We show that the time series of these three futures prices are cointegrated and we introduce a mean-reverting regime-switching process modulated by a hidden Markov chain to model the cointegration spread. By relying on this model and applying online filter-based parameter estimators, we implement and test several statistical arbitrage strategies. Our analysis reveals that statistical arbitrage strategies involving the recently introduced Shanghai futures are profitable even under conservative levels of transaction costs and over different time periods. Statistical arbitrage strategies involving only two of these three futures contracts or the three traditional crude oil futures (Brent, WTI, Dubai) deliver a lower investment performance.
@@ 1. Introduction
In this work, we study statistical arbitrage strategies in international crude oil futures markets, inspired by pairs trading strategies. The futures contracts we consider are two established benchmark crude oil futures, the Brent and the West Texas Intermediate (WTI), together with the Shanghai crude oil futures more recently introduced in 2018111The Shanghai International Energy Exchange (INE) introduced crude oil futures on March 26, 2018. These contracts are known as the Shanghai crude oil futures and are traded on the Shanghai Futures Exchange (SHFE). They are quoted in Chinese yuan and based on the INE’s own benchmark, which is a blend of various Middle Eastern and Asian crude oil grades that reflects the specific needs of China’s oil imports.. It known that the Brent and the WTI futures prices are cointegrated (see, e.g., [Kristoufek and Vosvrda, 2014], [Cerqueti et al., 2019], [Cerqueti and Fanelli, 2021], [Cotter et al., 2022]), while preliminary evidence (see [Yang and Zhou, 2020]) shows that the recently introduced Shanghai crude oil futures tends to cointegrate with them as well. These empirical findings suggest the possibility of achieving statistical arbitrage opportunities when these three crude oil futures are jointly traded. This represents the starting point and the motivation of this work.
We examine whether the introduction of the Shanghai crude oil futures has enabled international arbitrageurs to achieve profitable investments through statistical arbitrage strategies. Pairs trading on traditional crude oil benchmarks has already been studied (see, e.g., [Dunis et al., 2006], [Cummins and Bucca, 2012], [Baviera and Baldi, 2019]), while [Niu et al., 2023] test pairs trading strategies on the different Shanghai futures contracts traded at the INE. To the best of our knowledge, our work is the first to consider statistical arbitrage strategies involving simultaneously more than two futures contracts and, in particular, the established crude oil futures together with the recently introduced Shanghai crude oil futures.
@@
Our empirical analysis shows that statistical arbitrage strategies involving the three futures (Brent, WTI and Shanghai) at the same time, rather than just two of them, deliver significant investment performances even when transaction costs are accounted for. Moreover, strategies that exploit our hidden Markov model for the cointegration spread are remarkably profitable when compared to more traditional strategies based only on time series features. We find that the investment performances of these strategies are significantly reduced if the Shanghai futures is replaced by another standard crude oil benchmark as the Dubai crude oil futures. Our findings indicate that the greater profitability of statistical arbitrage strategies involving the Shanghai futures can be explained by the higher speed of adjustment of the Shanghai futures prices compared to the Brent and the WTI. This implies that the Shanghai futures prices tend to revert back quickly to the long-run relationship, making easier the exploitation of temporary mispricings. This finding is in line with the fact that the profitability of statistical arbitrage strategies tends to decline over time as markets develop (see, e.g., [Do and Faff, 2010]), while the introduction of a new security can represent a valuable opportunity for arbitrageurs [^2] See for instance [Crépellière et al., 2023] for an analysis of arbitrage opportunities in the emerging markets for cryptocurrencies..
@@ 2.1. A mean-reverting hidden Markov model
Let $F^{i}=(F^{i}_{t})_{t\geq 0}$ denote the futures price process [^3] We consider statistical arbitrage strategies based on the price processes of several assets. We refer to [Focardi et al., 2016] for an analysis of statistical arbitrage opportunities based on asset returns rather than prices., for $i\in\{B,S,W\}$, of the Brent ($B$), the Shanghai ($S$) and the WTI futures ($W$). The processes $F^{B}$ and $F^{W}$ are typically found to be cointegrated, as documented by [Hammoudeh et al., 2008] and [Kristoufek and Vosvrda, 2014] among others. Similarly, [Yang and Zhou, 2020] and [Huang and Huang, 2020] show that $F^{S}$ is cointegrated with $F^{B}$ and with $F^{W}$. Therefore, it is reasonable to guess that the three futures price processes jointly considered are cointegrated, similarly to the case of the Brent, the Dubai and the WTI futures prices (see [Galay, 2019]). The cointegration among $F^{B}$, $F^{S}$, $F^{W}$ has not been tested in previous works and will be shown to hold in the following. In the presence of cointegration, there exists a linear combination of $F^{B}$, $F^{S}$, $F^{W}$ that follows a stationary and possibly mean-reverting process. As usual in cointegration analysis, we call this linear combination the spread process and denote it by $S=(S_{t})_{t\geq 0}$. More specifically,
(2.1) $S_{t}:=\lambda^{0}+\sum_{i=B,S,W}\lambda^{i}F_{t}^{i},$
@@
Using the cointegration vector estimated in-sample an investor can evaluate the spread out-of-sample and set up statistical arbitrage strategies based on it, taking as portfolio weights the elements of the cointegration vector. These strategies turn out to be profitable as long as the cointegrating relationship is stable, namely as long as the spread $S$ is stationary. In a standard cointegration analysis, the cointegration vector $\lambda$ is assumed to be constant. However, previous works on the Brent and on the WTI document the existence of structural breaks in the cointegration vector (see, e.g., [Caporin et al., 2019]), showing also that the cointegration vector depends significantly on the time window over which the relationship is estimated. As already noticed by [Lee and Papanicolaou, 2016], this might severely impact pairs trading strategies based on cointegration. It is therefore appropriate to model the spread as a stochastic process with regime switching, updating the model parameters dynamically by incorporating new information as soon as it becomes available. In this work, we shall adopt this approach, starting from the explicit modelling of the spread process.
@@
Let $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space supporting a Brownian motion $W=(W_{t})_{t\geq 0}$ and a Markov chain $\textbf{X}=(\textbf{X}_{t})_{t\geq 0}$ with $N$ states and transition matrix $\boldsymbol{\Pi}$. The states of the Markov chain X represent different market regimes, thus capturing the time-varying nature of the cointegrating relationship. The Markov chain X admits a semimartingale representation of the form
$\textbf{X}_{t}=\textbf{X}_{0}+\int_{0}^{t}\boldsymbol{\Pi}\textbf{X}_{s}\mathrm{d}s+\textbf{M}_{t},$
@@
where $\textbf{M}=(\textbf{M}_{t})_{t\geq 0}$ is a martingale with $\textbf{M}_{0}=0$ (see [Elliott et al., 1995], Chapter 7).
@@
For the spread process $S$, we assume mean-reverting dynamics with regime switching:
(2.2) $\mathrm{d}S_{t}=a(\textbf{X}_{t})\left(\beta(\textbf{X}_{t})-S_{t}\right)\mathrm{d}t+\xi(\textbf{X}_{t})\mathrm{d}W_{t},\qquad\text{ with }S_{0}=s_{0}\in\mathbb{R},$
where $\alpha$, $\beta$, $\xi$ are real-valued functions on $\mathbb{R}^{N}$. Without loss of generality and similarly as in [Elliott et al., 1995], we assume that the state space of the Markov chain X is the canonical basis of $\mathbb{R}^{N}$, denoted by $\{\textbf{e}_{1},\ldots,\textbf{e}_{N}\}$, so that $a(\textbf{X}_{t})=\textbf{a}^{\top}\textbf{X}_{t}$ with $\textbf{a}\in\mathbb{R}^{N}$. For each $i=1,\dots,N$, the component $a_{i}$ represents the speed of mean-reversion of $S$ in the $i^{\text{th}}$ regime. Analogously, $\beta_{i}$ (resp. $\xi_{i}$) represents the long-run mean (resp. volatility) of $S$ in the $i^{\text{th}}$ regime. The resulting model for $S$ is a Ornstein–Uhlenbeck process driven by a hidden Markov model (OU-HMM henceforth) which has been used in the literature for the modelling of interest rates ([Erlwein and Mamon, 2009], [Grimm et al., 2020]), commodity spot prices ([Erlwein et al., 2010]) and spreads between equity price processes ([Elliott and Bradrania, 2018]).
@@
Since the current market regime is generally unknown by market participants, we assume that the Markov chain X is unobservable. As a consequence, the current state of X has to be filtered from the observations of the spread. Since we want to implement statistical arbitrage strategies that exploit the probabilistic structure of the model (2.2) for $S$, we need to estimate $\boldsymbol{a}$, $\boldsymbol{\beta}$, $\boldsymbol{\xi}$ and the transition matrix $\boldsymbol{\Pi}$. In this partial observation setup, the estimation of $\boldsymbol{a}$, $\boldsymbol{\beta}$, $\boldsymbol{\xi}$ and $\boldsymbol{\Pi}$ can be done by a filter-based Expectation Maximization algorithm that we now describe.
@@ 2.2. The filter-based Expectation Maximization algorithm
We now describe the filter-based Expectation Maximization (EM) algorithm for the estimation of the OU-HMM in (2.2). The following results are adapted from the general techniques described in [Elliott et al., 1995] and specifically applied to an Ornstein-Uhlenbeck process in [Erlwein and Mamon, 2009].
As a first step, we discretize the continuous-time model (2.2). This discretization is needed not only because only discrete observations of $S$ are available in practice, but also because in a continuous-time setting the EM algorithm does not allow to estimate $\boldsymbol{\xi}$, since it is based on equivalent measure transformations and volatility cannot be changed under equivalent measures.
Consider the spread process $S$ over a time interval $[t,t+\Delta]$, with $\Delta$ denoting the time step. If $\mathbf{X}$ is constant over this interval, we can write explicitly the solution of (2.2) as
$S_{t+\Delta}=e^{-a(\textbf{X}_{t})\Delta}S_{t}+\beta(\textbf{X}_{t})(1-e^{-a(\textbf{X}_{t})\Delta})+\xi(\textbf{X}_{t})\int_{t}^{t+\Delta}e^{-a(\textbf{X}_{t})(t+\Delta-s)}\mathrm{d}W_{s}.$
From this equation we derive the discrete-time version $y=(y_{t})_{t\in\mathbb{N}}$ of the spread:
(2.3) $y_{t+1}=\gamma(\textbf{X}_{t})+\alpha(\textbf{X}_{t})y_{t}+\eta(\textbf{X}_{t})z_{t},$
where $(z_{t})_{t\in\mathbb{N}}$ is a sequence of i.i.d. standard normal random variables and
(2.4) $\begin{array}[]{l}\alpha(\textbf{X}_{t}):=e^{-a(\textbf{X}_{t})\Delta},\\ \gamma(\textbf{X}_{t}):=\beta(\textbf{X}_{t})\left(1-e^{-a(\textbf{X}_{t})\Delta}\right),\\ \eta(\textbf{X}_{t}):=\xi(\textbf{X}_{t})\sqrt{\frac{1-e^{-2a(\textbf{X}_{t})\Delta}}{2a(\textbf{X}_{t})}}.\end{array}$
Analogously, we consider a discrete-time version of the Markov chain $\mathbf{X}$ as
$\mathbf{X}_{t+1}=\boldsymbol{\Pi}\mathbf{X}_{t}+\mathbf{v}_{t+1},$
with $\mathbf{v}_{t+1}$ denoting a martingale increment. We also introduce the following auxiliary quantities that will be needed for the recursive estimation of the model parameters:
@@
  • $T^{i}_{t}(f)=\sum_{n=1}^{t}\langle\mathbf{X}_{n-1},\mathbf{e}_{i}\rangle f_{n}$, where $f_{n}$ is a generic function of the observations of $y$ up to time $n$ (in our case, the function $f_{n}$ will be given by $f_{n}=y_{n}$, $f_{n}=y_{n}^{2}$ or $f_{n}=y_{n}y_{n-1}$).
We denote by $\mathbb{F}^{y}=(\mathcal{F}_{t}^{y})_{t\in\mathbb{N}}$, with $\mathcal{F}^{y}_{t}:=\sigma\left\{y_{0},\dots,y_{t}\right\}$ for all $t\in\mathbb{N}$, the filtration generated by the process $y$. We assume that $\mathbb{F}^{y}$ represents the information available to the investor, who cannot observe the Markov chain X and can only access discrete-time observations of the spread.
We now state the recursive filtering equations for the unobserved Markov chain $\mathbf{X}$ and the quantities $J$, $O$, $T$, from which we derive the recursive equations for the EM estimators of $\boldsymbol{\alpha}$, $\boldsymbol{\gamma}$, $\boldsymbol{\eta}$ and $\boldsymbol{\Pi}$. For $t\in\mathbb{N}$, we denote by $\hat{\mathbf{X}}_{t}:=\mathbb{E}\left[\mathbf{X}_{t}|\mathcal{F}^{y}_{t}\right]$ the filtered estimate of the latent Markov chain at time $t$ on the basis of the information generated by the observation process $y$ up to time $t$. In an analogous way we define the filtered estimates $\hat{J}_{t}^{ij}$, $\hat{O}_{t}^{i}$ and $\hat{T}_{t}^{i}(f)$ of the quantities introduced above. Moreover, we denote by $\hat{\boldsymbol{\alpha}}^{(t)}=(\hat{\alpha}_{i}^{(t)})_{i=1,\dots,N}$ the estimate of the parameter $\boldsymbol{\alpha}$ on the basis of the information generated by the observation process $y$ up to time $t$. The quantities $\hat{\boldsymbol{\gamma}}^{(t)}$, $\hat{\boldsymbol{\eta}}^{(t)}$, $\hat{\boldsymbol{\Pi}}^{(t)}$ are defined analogously.
As explained in [Elliott et al., 1995, Chapter 8], in order to filter the unobserved Markov chain $\mathbf{X}$ one can resort to a change of measure passing from $\mathbb{P}$ to an equivalent probability measure $\tilde{\mathbb{P}}$ under which the observation process $y$ is independent of the Markov chain $\mathbf{X}$. For each $T\in\mathbb{N}$, the probability measures $\mathbb{P}$ are $\tilde{\mathbb{P}}$ are related by the Radon-Nikodym derivative
@@
(2.5) $\lambda_{t}=\frac{1}{\eta(\mathbf{X}_{t-1})}\exp\left(-\frac{1}{2}\left(\frac{\left(y_{t}-y_{t-1}\alpha(\mathbf{X}_{t-1})-\gamma(\mathbf{X}_{t-1})\right)^{2}}{\eta(\mathbf{X}_{t-1})^{2}}-y_{t}^{2}\right)\right),$
@@
(2.6) $d^{ii}_{t}=\displaystyle\frac{1}{\eta_{i}}\exp\left(-\frac{1}{2}\left(\frac{\left(y_{t}-y_{t-1}\alpha_{i}-\gamma_{i}\right)^{2}}{\eta_{i}^{2}}-y_{t}^{2}\right)\right),\qquad\text{ for }i=1,\ldots,N.$
The starting point of the filtering algorithm is a set of initial guesses for the quantities $\mathbf{X}_{0}$ (that delivers also the initial values of the $O^{i}_{0}$’s) and for $\boldsymbol{\Pi}$, $\boldsymbol{\gamma},\boldsymbol{\alpha}$, $\boldsymbol{\eta}$.
@@
(2.8) $\begin{array}[]{rl}\hat{\pi}_{ij}^{(t+1)}&=\displaystyle\frac{\hat{J}^{i,j}_{t+1}}{\hat{O}^{j}_{t+1}},\\ \hat{\gamma}_{i}^{(t+1)}&=\displaystyle\frac{\hat{T}_{t+1}^{i}(y_{t+1})-\hat{\alpha}_{i}^{(t)}\hat{T}_{t}^{i}(y_{t})}{\hat{O}^{i}_{t+1}},\\ \hat{\alpha}_{i}^{(t+1)}&=\displaystyle\frac{\hat{T}_{t}^{i}(y_{t+1},y_{t})-\hat{\gamma}_{i}^{(t+1)}\hat{T}_{t}^{i}(y_{t})}{\hat{T}_{t}^{i}(y_{t}^{2})},\\ \hat{\eta}_{i}^{(t+1)}&=\displaystyle\frac{\hat{T}_{t+1}^{i}(y_{t+1}^{2})+(\hat{\alpha}_{i}^{(t+1)})^{2}\hat{T}_{t}^{i}(y_{t}^{2})+(\hat{\gamma}_{i}^{(t+1)})^{2}\hat{O}^{i}_{t+1}}{\hat{T}_{t}^{i}(y_{t}^{2})}\\ &\quad+\displaystyle\frac{\hat{\gamma}_{i}^{(t+1)}\hat{T}_{t+1}^{i}(y_{t+1})-2\hat{\alpha}_{i}^{(t+1)}\hat{T}_{t}^{i}(y_{t+1},y_{t})-2\hat{\alpha}_{i}^{(t+1)}\hat{\gamma}_{i}^{(t+1)}\hat{T}_{t}^{i}(y_{t})}{\hat{O}^{i}_{t+1}}.\\ \end{array}$
@@
In this remark we address some numerical aspects that arise when implementing the above filtering algorithm in practice:

### v3

@@ Abstract.
We study statistical arbitrage strategies in international crude oil futures markets. We analyse strategies that extend classical pairs trading strategies, considering two benchmark crude oil futures (Brent and WTI) together with the recently introduced Shanghai crude oil futures. We show that the time series of these three futures prices are cointegrated and we introduce a mean-reverting regime-switching process modulated by a hidden Markov chain to model the cointegration spread. By relying on this model and applying online filter-based parameter estimators, we implement and test several statistical arbitrage strategies. Our analysis shows that: (i) arbitrage strategies involving the recently introduced Shanghai futures can be profitable even under conservative levels of transaction costs; (ii) strategies based on our model outperform those relying solely on observed spread values; and (iii) incorporating three futures contracts enables the implementation of arbitrage strategies even in cases where pairwise cointegration is not detected.
@@ 1. Introduction
In this work, we study statistical arbitrage strategies in international crude oil futures markets, inspired by pairs trading strategies. The futures contracts we consider are two established benchmark crude oil futures, the Brent and the West Texas Intermediate (WTI), together with the Shanghai crude oil futures more recently introduced in 2018111The Shanghai International Energy Exchange (INE) introduced crude oil futures on March 26, 2018. These contracts are known as the Shanghai crude oil futures and are traded on the Shanghai Futures Exchange (SHFE). They are quoted in Chinese yuan and based on the INE’s own benchmark, which is a blend of various Middle Eastern and Asian crude oil grades that reflects the specific needs of China’s oil imports.. It is known that the Brent and the WTI futures prices are usually found to be cointegrated (see, e.g., [Kristoufek and Vosvrda, 2014], [Cerqueti et al., 2019], [Cerqueti and Fanelli, 2021], [Cotter et al., 2022]). Nevertheless, the Shanghai crude oil futures contract exhibits institutional and structural differences that can contribute to temporary segmentation from these traditional benchmarks. These differences include its denomination in Chinese Renminbi, trading restrictions and a benchmark composition based on Middle Eastern and Asian crude grades. Recent studies (e.g., [Yang et al., 2020], [Huang and Huang, 2020]) indicate that while the Shanghai market is gradually becoming more aligned with global benchmarks, it still shows signs of short-term divergences that present attractive opportunities for our statistical arbitrage strategies. These empirical findings motivate our joint cointegration analysis involving Brent, WTI, and Shanghai futures, and provide a compelling rationale for exploring potential statistical arbitrage opportunities in a globally integrated yet partially segmented crude oil futures market.
We examine whether the introduction of the Shanghai crude oil futures has enabled international arbitrageurs to achieve profitable investments through statistical arbitrage strategies. Pairs trading on traditional crude oil benchmarks has already been studied (see, e.g., [Dunis et al., 2006], [Cummins and Bucca, 2012], [Baviera and Baldi, 2019]), while [Niu et al., 2023] test pairs trading strategies on the different Shanghai futures contracts traded at the INE. To the best of our knowledge, our work is the first to consider statistical arbitrage strategies involving the recently introduced Shanghai crude oil futures.
@@
Our empirical analysis shows that statistical arbitrage strategies involving the three futures (Brent, WTI and Shanghai) at the same time deliver significant investment performances even when transaction costs are accounted for and even in cases where cointegration does not hold between any individual pair of these contracts. Moreover, strategies that exploit our hidden Markov model for the cointegration spread are remarkably profitable when compared to more traditional strategies based only on time series features. Our findings indicate that the greater profitability of statistical arbitrage strategies involving the Shanghai futures can be explained by the higher speed of adjustment of the Shanghai futures prices compared to the Brent and the WTI. This implies that the Shanghai futures prices tend to revert back quickly to the long-run relationship, making easier the exploitation of temporary mispricings. This finding is in line with the fact that the profitability of statistical arbitrage strategies tends to decline over time as markets develop (see, e.g., [Do and Faff, 2010]), while the introduction of a new security can represent a valuable opportunity for arbitrageurs [^2] See for instance [Crépellière et al., 2023] for an analysis of arbitrage opportunities in the emerging markets for cryptocurrencies..
@@ 2.1. A mean-reverting hidden Markov model
Let $F^{i}=(F^{i}_{t})_{t\in\mathbb{N}}$ denote the futures price process [^3] We consider statistical arbitrage strategies based on the price processes of several assets. We refer to [Focardi et al., 2016] for an analysis of statistical arbitrage opportunities based on asset returns rather than prices., for $i\in\{B,S,W\}$, of the Brent ($B$), the Shanghai ($S$), and the WTI futures ($W$). The processes $F^{B}$ and $F^{W}$ are typically found to be cointegrated, as documented by [Hammoudeh et al., 2008] and [Kristoufek and Vosvrda, 2014] among others. Similarly, [Yang and Zhou, 2020] and [Huang and Huang, 2020] show that $F^{S}$ is cointegrated with $F^{B}$ and with $F^{W}$. Therefore, it is reasonable to guess that the three futures price processes jointly considered are cointegrated, similarly to the case of the Brent, the Dubai and the WTI futures prices (see [Galay, 2019]). The cointegration among $F^{B}$, $F^{S}$, $F^{W}$ has not been tested in previous works and will be shown to hold in the following. In the presence of cointegration, there exists a linear combination of $F^{B}$, $F^{S}$, $F^{W}$ that follows a stationary and possibly mean-reverting process. As usual in cointegration analysis, we call this linear combination the spread process and denote it by $S=(S_{t})_{t\in\mathbb{N}}$:
(2.1) $S_{t}:=\lambda^{0}+\sum_{i=B,S,W}\lambda^{i}F_{t}^{i},\qquad\text{ for all }t\in\mathbb{N},$
@@
Using the cointegration vector estimated in-sample, an investor can evaluate the spread out-of-sample and set up statistical arbitrage strategies based on it, taking as portfolio weights the elements of the cointegration vector (see Section 3 for more details). These strategies turn out to be profitable as long as the cointegrating relationship is stable, namely as long as the spread $S$ is stationary. In a standard cointegration analysis, the cointegration vector $\lambda$ is assumed to be constant over time. However, previous works on the Brent and on the WTI document the presence of structural breaks in the cointegration vector (see, e.g., [Caporin et al., 2019]), showing also that the cointegration vector depends significantly on the time window under consideration. As already noted by [Lee and Papanicolaou, 2016], this might severely impact the performance of pairs trading strategies based on cointegration. It is therefore appropriate to model the spread as a stochastic process with regime switching, updating dynamically the model parameters as new information becomes available. In this work, we adopt this approach and explicitly model the spread as a regime-switching stochastic process.
@@
Let $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{t\in\mathbb{N}},\mathbb{P})$ be a filtered probability space supporting a sequence $(z_{t})_{t\in\mathbb{N}}$ of i.i.d. standard normal random variables and a Markov chain $\textbf{X}=(\textbf{X}_{t})_{t\in\mathbb{N}}$ with $N$ states and transition matrix $\boldsymbol{\Pi}$. The states of the Markov chain X represent different market regimes, thus capturing the time-varying nature of the cointegrating relationship. The Markov chain X admits a representation of the form
$\textbf{X}_{t+1}=\boldsymbol{\Pi}\textbf{X}_{t}+\textbf{M}_{t+1},\qquad\text{ for all }t\in\mathbb{N},$
@@
where $\textbf{X}_{0}$ is the initial state of the Markov chain and $\textbf{M}=(\textbf{M}_{t})_{t\in\mathbb{N}}$ is a sequence of martingale increments with $\textbf{M}_{0}=0$ (see [Elliott et al., 1995], Chapter 2).
@@
For the spread process $S$, we assume autoregressive dynamics with regime switching:
(2.2) $S_{t+1}=\gamma(\textbf{X}_{t})+\alpha(\textbf{X}_{t})S_{t}+\eta(\textbf{X}_{t})z_{t},\qquad S_{0}=s_{0}\in\mathbb{R},$
where $\gamma$, $\alpha$ and $\eta$ are real-valued functions on $\mathbb{R}^{N}$. Without loss of generality and similarly as in [Elliott et al., 1995], we assume that the state space of the Markov chain X is the canonical basis of $\mathbb{R}^{N}$, denoted by $\{\textbf{e}_{1},\ldots,\textbf{e}_{N}\}$, so that $\gamma(\textbf{X}_{t})=\boldsymbol{\gamma}^{\top}\textbf{X}_{t}$ with $\boldsymbol{\gamma}\in\mathbb{R}^{N}$. The term $\gamma_{i}$ represents the intercept of $S$ in the $i^{\text{th}}$ regime, while $\alpha_{i}$ and $\eta_{i}$ represent respectively the autoregressive coefficient and the volatility of $S$ in the $i^{\text{th}}$ regime, for each $i=1,\dots,N$. The condition $|\alpha_{i}|<1$, ensures that the spread is mean-reverting within each regime and admits a unique stationary distribution conditionally on that regime. The resulting model for $S$ is an autoregressive process of order one modulated by a Markov chain, widely used in the time-series literature to capture nonlinear dynamics and regime-dependent persistence (see, e.g., [Hamilton, 1994], Chapter 22), which can effectively affect the spread process in our context.
@@
Since the current market regime is generally unknown by market participants, we assume that the Markov chain X is unobservable, giving rise to an autoregressive process modulated by a hidden Markov model (AR-HMM). As a consequence, the current state of X has to be filtered from the observations of the spread. Since we want to implement statistical arbitrage strategies that exploit the probabilistic structure of the model (2.2) for $S$, we need to estimate $\boldsymbol{\gamma}$, $\boldsymbol{\alpha}$, $\boldsymbol{\eta}$ and the transition matrix $\boldsymbol{\Pi}$. In this partial observation setup, the estimation of $\boldsymbol{\gamma}$, $\boldsymbol{\alpha}$, $\boldsymbol{\eta}$ and $\boldsymbol{\Pi}$ can be done by a filter-based Expectation Maximization algorithm that we describe below.
Remark 2.1.
The regime-switching autoregressive model introduced above can be seen as a discrete-time representation, at the observation frequency, of a continuous-time Ornstein-Uhlenbeck process modulated by a hidden Markov model (OU-HMM). Specifically, let $\textbf{X}=(\textbf{X}_{t})_{t\geq 0}$ be a continuous-time Markov chain. The continuous-time analogue of the discrete-time spread dynamics (2.2) can then be written as follows:
(2.3) $\mathrm{d}S_{t}=a(\textbf{X}_{t})\left(\beta(\textbf{X}_{t})-S_{t}\right)\mathrm{d}t+\xi(\textbf{X}_{t})\mathrm{d}W_{t},$
where $W=(W_{t})_{t\geq 0}$ is a Brownian motion. The functions $a$, $\beta$ and $\xi$ represent respectively the speed of mean-reversion, the long-run mean and the volatility of the spread as functions of the current regime. OU-HMMs have been used for the modelling of interest rates ([Erlwein and Mamon, 2009], [Grimm et al., 2020]), commodity spot prices ([Erlwein et al., 2010]) and spreads between equity price processes ([Elliott and Bradrania, 2018]). The coefficients of (2.3) can be mapped into the coefficients of (2.2) as follows, where $\Delta$ denotes the observation frequency:
$\alpha(\textbf{X}_{t})=e^{-a(\textbf{X}_{t})\Delta},\qquad\gamma(\textbf{X}_{t})=\beta(\textbf{X}_{t})\bigl(1-e^{-a(\textbf{X}_{t})\Delta}\bigr),\qquad\eta(\textbf{X}_{t})=\xi(\textbf{X}_{t})\sqrt{\frac{1-e^{-2a(\textbf{X}_{t})\Delta}}{2a(\textbf{X}_{t})}}.$
While continuous-time mean-reverting dynamics have been considered in several works on pairs trading (see for instance [Avellaneda and Lee, 2010, Bertram, 2010, Yun and He, 2022]), we emphasize that adopting the discrete-time formulation in (2.2) is not only motivated by the fact that observations are available at discrete intervals, but is also essential from an estimation perspective. In continuous time, the EM algorithm relies on equivalent measure transformations, under which volatility parameters are invariant. As a consequence, the volatility function cannot be estimated by a continuous-time EM algorithm, whereas the discrete-time formulation allows all model parameters, including volatility, to be consistently estimated.
@@ 2.2. The filter-based Expectation Maximization algorithm
We now describe the filter-based Expectation Maximization (EM) algorithm for the estimation of the AR-HMM in (2.2). The following results are adapted from the general techniques described in [Elliott et al., 1995] and more specifically in [Erlwein and Mamon, 2009] for HMM-modulated autoregressive processes with Gaussian innovations. As a preliminary, we introduce the following auxiliary quantities that will be needed for the recursive estimation of the model parameters:
@@
  • $T^{i}_{t}(f)=\sum_{n=1}^{t}\langle\mathbf{X}_{n-1},\mathbf{e}_{i}\rangle f_{n}$, where $f_{n}$ is a generic function of the observations of $S$ up to time $n$ (in our case, the function $f_{n}$ will be given by $f_{n}=S_{n}$, $f_{n}=S_{n}^{2}$ or $f_{n}=S_{n}S_{n-1}$).
We denote by $\mathbb{F}^{S}=(\mathcal{F}_{t}^{S})_{t\in\mathbb{N}}$, with $\mathcal{F}^{S}_{t}:=\sigma\left\{S_{0},\dots,S_{t}\right\}$ for all $t\in\mathbb{N}$, the filtration generated by the spread process $S$. We assume that $\mathbb{F}^{S}$ represents the information available to the investor, who cannot observe the Markov chain X and can only access current and past discrete-time observations of the spread.
We now state the recursive filtering equations for the unobserved Markov chain $\mathbf{X}$ and the quantities $J$, $O$, $T$, from which we derive the recursive equations for the EM estimators of $\boldsymbol{\alpha}$, $\boldsymbol{\gamma}$, $\boldsymbol{\eta}$ and $\boldsymbol{\Pi}$. Those filtered estimates will enable the investor to compute one-step-ahead forecasts of the conditional mean and variance of the spread, which will be used to design model-based statistical arbitrage strategies in Section 3. For $t\in\mathbb{N}$, we denote by $\hat{\mathbf{X}}_{t}:=\mathbb{E}\left[\mathbf{X}_{t}|\mathcal{F}^{S}_{t}\right]$ the filtered estimate of the latent Markov chain at time $t$ on the basis of the information generated by the observation process $S$ up to time $t$. In an analogous way we define the filtered estimates $\hat{J}_{t}^{ij}$, $\hat{O}_{t}^{i}$ and $\hat{T}_{t}^{i}(f)$ of the quantities introduced above. Moreover, we denote by $\hat{\boldsymbol{\alpha}}^{(t)}=(\hat{\alpha}_{i}^{(t)})_{i=1,\dots,N}$ the estimate of the parameter $\boldsymbol{\alpha}$ on the basis of the information generated by the observation process $S$ up to time $t$. The quantities $\hat{\boldsymbol{\gamma}}^{(t)}$, $\hat{\boldsymbol{\eta}}^{(t)}$, $\hat{\boldsymbol{\Pi}}^{(t)}$ are defined analogously.
As explained in [Elliott et al., 1995, Chapter 2], in order to filter the unobserved Markov chain $\mathbf{X}$ one can resort to a change of measure passing from $\mathbb{P}$ to an equivalent probability measure $\tilde{\mathbb{P}}$ under which the observation process $S$ is independent of the Markov chain $\mathbf{X}$. For each $T\in\mathbb{N}$, the probability measures $\mathbb{P}$ are $\tilde{\mathbb{P}}$ are related by the Radon-Nikodym derivative
@@
(2.4) $\lambda_{t}=\exp\left(-\frac{1}{\eta(\mathbf{X}_{t-1})^{2}}\left(\left(\alpha(\mathbf{X}_{t-1})S_{t-1}+\gamma(\mathbf{X}_{t-1})\right)S_{t}+\frac{1}{2}\left(\alpha(\mathbf{X}_{t-1})S_{t-1}+\gamma(\mathbf{X}_{t-1})\right)^{2}\right)\right).$
@@
(2.5) $d^{ii}_{t}=\exp\left(-\frac{1}{\eta_{i}^{2}}\left(\left(\alpha_{i}S_{t-1}+\gamma_{i}\right)S_{t}+\frac{1}{2}\left(\alpha_{i}S_{t-1}+\gamma_{i}\right)^{2}\right)\right),\qquad\text{ for }i=1,\ldots,N.$
The starting point of the filtering algorithm is a set of initial guesses for the quantities $\mathbf{X}_{0}$ (that delivers also the initial values of $O^{i}_{0}$) and for $\boldsymbol{\Pi}$, $\boldsymbol{\gamma},\boldsymbol{\alpha}$, $\boldsymbol{\eta}$, see Remark 2.2 for more details.
@@
[…]
@@
In this remark we address some numerical aspects that have to be taken into account when implementing the above filtering algorithm in practice: